OpenStudy (anonymous):

How do I find the inverse of f(x)= 3x+15 over -17x-16

OpenStudy (amistre64):

if it has an inverse; just solve it for the "x" variable

OpenStudy (amistre64):

y = f(x) = 3x+15 / -17x-16 y = 3x+15 / -17x-16 y(-17x -16) = 3x+15 -17xy -16y = 3x+15 -17xy -3x = 16y + 15 x(-17y -3) = 16y + 15 x = f(y) = 16y+15 / -17y-3